### v2

# Kaon-induced production of strange hidden-charm molecular pentaquarks $P^{\Lambda}_{\psi s}$ from proton

Abstract
In this work, we investigate the kaon-induced production of strange hidden-charm pentaquark $P^{\Lambda}_{\psi s}$ states, including the $P^{\Lambda}_{\psi s}(4459)$ and $P^{\Lambda}_{\psi s}(4338)$ observed by the LHCb collaboration. Coupled-channel interactions involving the channels $\bar{K}N$, $\Xi_{c}^{(*)}\bar{D}^{(*)}$, $\Xi^{\prime}_{c}\bar{D}^{(*)}$, $\Lambda_{c}\bar{D}_{s}^{(*)}$, and $\Lambda J/\psi$ are studied using effective Lagrangians to describe these interactions. The potential kernels are constructed within the one-boson-exchange model and implemented into the quasipotential Bethe-Salpeter equation to calculate the scattering amplitudes. From these amplitudes, the partial-wave cross sections for kaon-induced production are predicted alongside the poles corresponding to strange hidden-charm molecular pentaquarks. The analysis reveals complex structures in the cross sections due to molecular states. A structure near the $\Xi_{c}\bar{D}$ threshold is predicted in the $\Lambda_{c}\bar{D}_{s}$ channel, corresponding to $P^{\Lambda}_{\psi s}(4338)$. Both molecular states near the $\Xi_{c}\bar{D}^{*}$ threshold with $J^{P}=1/2^{-}$ and $J^{P}=3/2^{-}$, associated with $P^{\Lambda}_{\psi s}(4459)$, produce peaks in the $\Xi^{\prime}_{c}\bar{D}$ channel, which vary significantly with the cutoff. Additionally, a molecular state with $J^{P}=3/2^{-}$ near the $\Xi_{c}^{*}\bar{D}$ threshold gives rise to a peak in the $\Xi_{c}\bar{D}^{*}$ channel. In the higher energy region, from 4350 to 4550 MeV, although the cross sections in the $\Lambda_{c}\bar{D}_{s}^{(*)}$ channels are substantial, no distinct structures corresponding to strange hidden-charm pentaquarks are identified, apart from some threshold-related cusps. These results provide valuable theoretical insights and serve as a foundation for future experimental studies of strange hidden-charm pentaquarks using kaon beams at J-PARC and JLab.
journal: Eur. Phys. J. C
1 INTRODUCTION
With the progress of experimental studies, particularly the observations at LHCb, hidden-charm pentaquark states have garnered significant attention and inspired extensive global research. The story began with the discovery of the $P^{N}_{\psi}(4450)$ and $P^{N}_{\psi}(4380)$ states by the LHCb collaboration in 2015 Aaij:2015tga , confirming predictions made by several theorists Wu:2010jy ; Yang:2011wz ; Wang:2011rga ; Xiao:2013yca . This experimental breakthrough led to a surge of theoretical investigations into hidden-charm pentaquark structures Chen:2016qju ; Guo:2017jvc ; Chen:2015loa ; Chen:2015moa ; Karliner:2015ina ; Roca:2015dva ; He:2015cea ; Burns:2015dwa . The narrow widths and masses of these states, which are close to the corresponding thresholds, strongly support the molecular state interpretation. Specifically, $P^{N}_{\psi}(4380)$ and $P^{N}_{\psi}(4450)$ have been assigned as $\bar{D}\Sigma^{*}_{c}$ and $\bar{D}^{*}\Sigma_{c}$ molecular states, respectively, in our previous work He:2015cea . In subsequent observations, the $P^{N}_{\psi}(4450)$ state was resolved into two separate structures, $P^{N}_{\psi}(4440)$ and $P^{N}_{\psi}(4457)$, and a new state, $P^{N}_{\psi}(4312)$, was discovered in 2019 Aaij:2019vzc . These states are interpreted as S-wave molecular states composed of anticharmed mesons and charmed baryons, with their masses closely aligned with the relevant thresholds Liu:2019tjn ; Chen:2019asm ; Xiao:2019aya ; He:2019ify ; Chen:2019bip ; Du:2019pij . Given the success of the molecular picture in describing non-strange hidden-charm systems, it is natural to extend this framework to hidden-charm pentaquark states with strangeness Anisovich:2015zqa ; Wang:2015wsa ; Feijoo:2015kts ; Lu:2016roh ; Chen:2015sxa ; Chen:2016ryt ; Xiao:2019gjd ; Zhang:2020cdi ; Wang:2019nvm . The experimental observation of strange hidden-charm pentaquarks is therefore pivotal in advancing our understanding of molecular states with hidden charm.
In 2020, the LHCb Collaboration reported the observation of the strange hidden-charm pentaquark $P^{\Lambda}_{\psi s}(4459)$ in the $\Xi_{b}^{-}\to J/\psi\Lambda K^{-}$ decay Aaij:2020gdg . This state has a measured mass approximately 19 MeV below the $\Xi_{c}\bar{D}^{*}$ threshold and a width of 17 MeV Aaij:2020gdg , characteristics consistent with its interpretation as a molecular state. Numerous studies have investigated the molecular state nature of $P^{\Lambda}_{\psi s}(4459)$. For instance, our previous work suggests that the structure of $P^{\Lambda}_{\psi s}(4459)$ primarily arises from the $\Xi_{c}\bar{D}^{*}$ configuration with $J^{P}=1/2^{-}$, while the possible contribution of the $\Xi_{c}\bar{D}^{*}$ state with $J^{P}=3/2^{-}$ cannot be excluded Zhu:2021lhd ; Zhu:2022wpi . Similarly, QCD sum rule analyses in Ref. Chen:2020uif support the interpretation of $P^{\Lambda}_{\psi s}(4459)$ as a $\Xi_{c}\bar{D}^{*}$ molecular state with $J^{P}=1/2^{-}$ or $3/2^{-}$. Ref. Wang:2021itn interprets $P^{\Lambda}_{\psi s}(4459)$ as either a $\Xi^{\prime}_{c}\bar{D}$ molecular state with quantum numbers $I(J^{P})=0(3/2^{-})$ or a $\Xi_{c}\bar{D}^{*}$ molecular state with $I(J^{P})=0(3/2^{-})$. Moreover, studies within the framework of heavy quark spin symmetry (HQSS), such as those in Ref. Peng:2020hql , favor the $\Xi_{c}\bar{D}^{*}$ configuration with $J^{P}=3/2^{-}$ as the most likely candidate for $P^{\Lambda}_{\psi s}(4459)$, ruling out the $J^{P}=1/2^{-}$ possibility. A recent study suggests that the strong coupling between $\bar{D}_{s}\Lambda_{c}$ and $\bar{D}\Xi_{c}$, as well as between $\bar{D}s^{*}\Lambda_{c}$ and $\bar{D}^{*}\Xi_{c}$, leads to the association of $P^{\Lambda}_{\psi s}(4338)$ and $P^{\Lambda}_{\psi s}(4459)$ with states predominantly coupling to $\bar{D}\Xi^{\prime}$ and $\bar{D}^{*}\Xi_{c}$, respectively Feijoo:2022rxf .
In 2022, the LHCb Collaboration identified a new neutral strange hidden-charm pentaquark state, $P^{\Lambda}_{\psi s}(4338)$, in the $B^{-}\to J/\psi\Lambda\bar{p}$ decay LHCb:2022ogu . The state was measured to have a mass of $4338.3\pm 0.7\pm 0.4$ MeV and a width of $7.0\pm 1.2\pm 1.3$ MeV. The mass and narrow width of $P^{\Lambda}_{\psi s}(4338)$ are consistent with the characteristics of a molecular state, which has inspired several studies Wang:2022neq ; Wang:2022mxy ; Yan:2022wuz ; Ozdem:2022kei ; Wang:2022tib ; Nakamura:2022gtu ; Giachino:2022pws ; Ortega:2022uyu proposing its molecular nature. In our previous work, the lineshape of the $P^{\Lambda}_{\psi s}(4338)$ structure can be approximately reproduced by a narrow molecular state resulting from the $\Xi_{c}\bar{D}$ interaction with $J^{P}=1/2^{-}$, which lies very close to the threshold and exhibits a significant interference effect Zhu:2022wpi .
Although hidden-charm pentaquarks, including those with strangeness, have been observed at LHCb, further confirmation and a deeper understanding of these states require analyzing their production mechanisms across different reaction channels. Even before the observation of the $P^{N}_{\psi}$ state at LHCb, the production of these states via proton-antiproton collisions and photoproduction was proposed Wu:2010jy ; He:2015cea ; Huang:2013mua . Following the LHCb observations, many studies were proposed to investigate the production of hidden-charm pentaquarks, $P^{\Lambda}_{\psi}$, via pion-induced reactions or photoproduction Lu:2015fva ; Huang:2016tcr ; Kim:2016cxr ; Cao:2019gqo ; Wang:2019krd ; Xie:2020niw ; Shi:2022ipx ; Wang:2019dsi . Additionally, the pion-induced production of strange hidden-charm pentaquarks was studied even before the LHCb observation Wang:2015xwa . In Refs. Clymton:2021thh ; Paryev:2023icm ; Paryev:2022zdx ; Clymton:2022qlr , the authors employed two different methods, the effective Lagrangian approach and the Regge method, to study the production of the exotic hidden-charm pentaquark state $P^{\Lambda}_{\psi s}(4459)$ in the $\bar{K}p\rightarrow J/\psi\Lambda$ reaction. In Ref. Cheng:2021gca , the authors explored the possibility of investigating the strange hidden-charm pentaquark state $P^{\Lambda}_{\psi s}(4459)$ through photon-induced reactions on a proton target in the $\gamma p\rightarrow K^{+}P^{\Lambda}_{\psi s}(4459)$ channel, using an effective Lagrangian approach.
In most studies of photoproduction or meson production of hidden-charm pentaquarks, information about the pentaquarks is extracted directly from experiments, and coupled-channel effects are not considered. In the current work, we aim to provide a more comprehensive study of the production of hidden-charm pentaquarks in molecular state picture through kaon-nucleon scattering. Building on our previous research Zhu:2021lhd ; Zhu:2022wpi , we calculate the coupled-channel interactions between charmed baryons and anticharmed mesons within the framework of the one-boson-exchange model. In addition to the channels in the charmed energy region, including $\bar{K}N$, $\Xi_{c}^{(*)}\bar{D}^{(*)}$, $\Xi^{\prime}_{c}\bar{D}^{(*)}$, $\Lambda_{c}\bar{D}_{s}^{(*)}$, and $\Lambda J/\psi$, which successfully reproduce the experimentally observed $P^{\Lambda}_{\psi s}(4459)$ and $P^{\Lambda}_{\psi s}(4438)$ Zhu:2022wpi , we extend our study by incorporating the light channel $\bar{K}N$ for kaon-induced production. Using the quasipotential Bethe-Salpeter equation (qBSE), we calculate the scattering amplitudes, which can then be used to determine the cross sections for the transition from the $\bar{K}N$ channel to those in the charmed energy region.
This article is organized as follows. After the introduction, Sec. 2 outlines the Lagrangians used to construct the potentials for coupled-channel interactions, the qBSE approach, and the methodology for calculating the scattering cross sections. In Sec. 3, we present the results for the cross sections of kaon-induced production. Finally, Sec. 4 provides a summary of the entire work.
2 Formalism
In the present work, we calculate the cross sections by using coupled-channel interactions to obtain the scattering amplitudes for the $\bar{K}N$ and charmed channels within the quasipotential Bethe-Salpeter equation (qBSE) approach. To start, the potential kernel is derived from effective Lagrangians within the one-boson-exchange model.
2.1 Relevant lagrangians and potential kernels
In the present work, we consider molecular pentaquarks produced from the interaction of charmed baryons and anticharmed mesons, including the channels $\Xi_{c}^{*}\bar{D}^{*}$, $\Xi^{\prime}_{c}\bar{D}^{*}$, $\Xi^{*}_{c}\bar{D}$, $\Xi_{c}\bar{D}^{*}$, $\Xi^{\prime}_{c}\bar{D}$, $\Lambda_{c}\bar{D}_{s}$, $\Xi_{c}\bar{D}$, and $\Lambda_{c}\bar{D}_{s}$, with light meson exchange within the one-boson-exchange model. We also consider the channel involving a hidden-charm meson, $\Lambda J/\psi$, for which the interaction is neglected. Therefore, we require vertices for the coupling between the constituent particles and the exchanged meson. The relevant Lagrangians for these vertices are given as follows Zhu:2021lhd ; Cheng:1992xi ; Yan:1992gz ; Wise:1992hn ; Casalbuoni:1996pg ,
$\mathcal{L}_{\mathcal{\tilde{P}}^{*}\mathcal{\tilde{P}}\mathbb{P}}$ $=i\frac{2g\sqrt{m_{\mathcal{\tilde{P}}}m_{\mathcal{\tilde{P}}^{*}}}}{f_{\pi}}(-\mathcal{\tilde{P}}^{*{\dagger}}_{a\lambda}\mathcal{\tilde{P}}_{b}+\mathcal{\tilde{P}}^{\dagger}_{a}\mathcal{\tilde{P}}^{*}_{b\lambda})\partial^{\lambda}\mathbb{P}_{ab},$
$\mathcal{L}_{\mathcal{\tilde{P}}^{*}\mathcal{\tilde{P}}^{*}\mathbb{P}}$ $=-\frac{g}{f_{\pi}}\epsilon_{\alpha\mu\nu\lambda}\mathcal{\tilde{P}}^{*\mu{\dagger}}_{a}\overleftrightarrow{\partial}^{\alpha}\mathcal{\tilde{P}}^{*\lambda}_{b}\partial^{\nu}\mathbb{P}_{ba},$
$\mathcal{L}_{\mathcal{\tilde{P}}^{*}\mathcal{\tilde{P}}\mathbb{V}}$ $=\sqrt{2}\lambda g_{V}\varepsilon_{\lambda\alpha\beta\mu}(-\mathcal{\tilde{P}}^{*\mu{\dagger}}_{a}\overleftrightarrow{\partial}^{\lambda}\mathcal{\tilde{P}}_{b}+\mathcal{\tilde{P}}^{\dagger}_{a}\overleftrightarrow{\partial}^{\lambda}\mathcal{\tilde{P}}_{b}^{*\mu})\partial^{\alpha}\mathbb{V}^{\beta}_{ab},$
$\mathcal{L}_{\mathcal{\tilde{P}}\mathcal{\tilde{P}}\mathbb{V}}$ $=i\frac{\beta g_{V}}{\sqrt{2}}\mathcal{\tilde{P}}_{a}^{\dagger}\overleftrightarrow{\partial}_{\mu}\mathcal{\tilde{P}}_{b}\mathbb{V}^{\mu}_{ab},$
$\mathcal{L}_{\mathcal{\tilde{P}}^{*}\mathcal{\tilde{P}}^{*}\mathbb{V}}$ $=-i\frac{\beta g_{V}}{\sqrt{2}}\mathcal{\tilde{P}}_{a}^{*{\dagger}}\overleftrightarrow{\partial}_{\mu}\mathcal{\tilde{P}}^{*}_{b}\mathbb{V}^{\mu}_{ab}$
$-i2\sqrt{2}\lambda g_{V}m_{\mathcal{\tilde{P}}^{*}}\mathcal{\tilde{P}}^{*\mu{\dagger}}_{a}\mathcal{\tilde{P}}^{*\nu}_{b}(\partial_{\mu}\mathbb{V}_{\nu}-\partial_{\nu}\mathbb{V}_{\mu})_{ab},$
$\mathcal{L}_{\mathcal{\tilde{P}}\mathcal{\tilde{P}}\sigma}$ $=-2g_{s}m_{\mathcal{\tilde{P}}}\mathcal{\tilde{P}}_{a}^{\dagger}\mathcal{\tilde{P}}_{a}\sigma,$
$\mathcal{L}_{\mathcal{\tilde{P}}^{*}\mathcal{\tilde{P}}^{*}\sigma}$ $=2g_{s}m_{\mathcal{\tilde{P}}^{*}}\mathcal{\tilde{P}}_{a}^{*{\dagger}}\mathcal{\tilde{P}}^{*}_{a}\sigma,$ (1)
where the $\mathcal{\tilde{P}}=(\bar{D}^{0},D^{-},D^{-}_{s})$, and the $\mathbb{P}$ and $\mathbb{V}$ are the pseudoscalar and vector matrices as
${\mathbb{P}}$ $=\left(\begin{array}[]{ccc}\frac{\sqrt{3}\pi^{0}+\eta}{\sqrt{6}}&\pi^{+}&K^{+}\\ \pi^{-}&\frac{-\sqrt{3}\pi^{0}+\eta}{\sqrt{6}}&K^{0}\\ K^{-}&\bar{K}^{0}&-\frac{2\eta}{\sqrt{6}}\end{array}\right),\mathbb{V}$ $=\left(\begin{array}[]{ccc}\frac{\rho^{0}+\omega}{\sqrt{2}}&\rho^{+}&K^{*+}\\ \rho^{-}&\frac{-\rho^{0}+\omega}{\sqrt{2}}&K^{*0}\\ K^{*-}&\bar{K}^{*0}&\phi\end{array}\right),$ (8)
where the indices $a,b=1,2,3$ are used to label the particle elements in the matrices ${\mathbb{P}},{\mathbb{V}}$ and vector $\mathcal{P}$.
For the baryon side, the Lagrangians for the couplings between the charmed baryons and light mesons are given as follows,
${\cal L}_{BB\mathbb{P}}$ $=-i\frac{3g_{1}}{4f_{\pi}\sqrt{m_{\bar{B}}m_{B}}}~{}\epsilon^{\mu\nu\lambda\kappa}\partial_{\nu}\mathbb{P}~{}\sum_{i,j=0,1}\bar{B}_{i\mu}\overleftrightarrow{\partial}_{\kappa}B_{j\lambda},$
${\cal L}_{BB\mathbb{V}}$ $=-\frac{\beta_{S}g_{V}}{2\sqrt{2m_{\bar{B}}m_{B}}}\mathbb{V}^{\nu}\sum_{i,j=0,1}\bar{B}_{i}^{\mu}\overleftrightarrow{\partial}_{\nu}B_{j\mu}$
$-\frac{\lambda_{S}g_{V}}{\sqrt{2}}(\partial_{\mu}\mathbb{V}_{\nu}-\partial_{\nu}\mathbb{V}_{\mu})\sum_{i,j=0,1}\bar{B}_{i}^{\mu}B_{j}^{\nu},$
${\cal L}_{BB\sigma}$ $=\ell_{S}\sigma\sum_{i,j=0,1}\bar{B}_{i}^{\mu}B_{j\mu},$
${\cal L}_{B_{\bar{3}}B_{\bar{3}}\mathbb{V}}$ $=-\frac{g_{V}\beta_{B}}{2\sqrt{2m_{\bar{B}_{\bar{3}}}m_{B_{\bar{3}}}}}\mathbb{V}^{\mu}\bar{B}_{\bar{3}}\overleftrightarrow{\partial}_{\mu}B_{\bar{3}},$
${\cal L}_{B_{\bar{3}}B_{\bar{3}}\sigma}$ $=i\ell_{B}\sigma\bar{B}_{\bar{3}}B_{\bar{3}},$
${\cal L}_{BB_{\bar{3}}\mathbb{P}}$ $=-i\frac{g_{4}}{f_{\pi}}\sum_{i}\bar{B}_{i}^{\mu}\partial_{\mu}\mathbb{P}B_{\bar{3}}+{\rm H.c.},$
${\cal L}_{BB_{\bar{3}}\mathbb{V}}$ $=\frac{g_{\mathbb{V}}\lambda_{I}}{\sqrt{2m_{\bar{B}}m_{B_{\bar{3}}}}}\epsilon^{\mu\nu\lambda\kappa}\partial_{\lambda}\mathbb{V}_{\kappa}\sum_{i}\bar{B}_{i\nu}\overleftrightarrow{\partial}_{\mu}B_{\bar{3}}+{\rm H.c.},$ (9)
where the Dirac spinor operators with label $i,j=0,1$ are defined as,
${B}_{0\mu}$ $\equiv-\sqrt{\frac{1}{3}}(\gamma_{\mu}+v_{\mu})\gamma^{5}B;\,\,\,B_{1\mu}\equiv B^{*}_{\mu},$
${\bar{B}}_{0\mu}$ $\equiv\sqrt{\frac{1}{3}}\bar{B}\gamma^{5}(\gamma_{\mu}+v_{\mu});\,\,\,\bar{B}_{1\mu}\equiv\bar{B}^{*}_{\mu},$ (10)
and the charmed baryon matrices are defined as
$B_{\bar{3}}$ $=\left(\begin{array}[]{ccc}0&\Lambda^{+}_{c}&\Xi_{c}^{+}\\ -\Lambda_{c}^{+}&0&\Xi_{c}^{0}\\ -\Xi^{+}_{c}&-\Xi_{c}^{0}&0\end{array}\right),\quad B=\left(\begin{array}[]{ccc}\Sigma_{c}^{++}&\frac{1}{\sqrt{2}}\Sigma^{+}_{c}&\frac{1}{\sqrt{2}}\Xi^{\prime+}_{c}\\ \frac{1}{\sqrt{2}}\Sigma^{+}_{c}&\Sigma_{c}^{0}&\frac{1}{\sqrt{2}}\Xi^{\prime 0}_{c}\\ \frac{1}{\sqrt{2}}\Xi^{\prime+}_{c}&\frac{1}{\sqrt{2}}\Xi^{\prime 0}_{c}&\Omega^{0}_{c}\end{array}\right),$ (17)
$B^{*}$ $=\left(\begin{array}[]{ccc}\Sigma_{c}^{*++}&\frac{1}{\sqrt{2}}\Sigma^{*+}_{c}&\frac{1}{\sqrt{2}}\Xi^{*+}_{c}\\ \frac{1}{\sqrt{2}}\Sigma^{*+}_{c}&\Sigma_{c}^{*0}&\frac{1}{\sqrt{2}}\Xi^{*0}_{c}\\ \frac{1}{\sqrt{2}}\Xi^{*+}_{c}&\frac{1}{\sqrt{2}}\Xi^{*0}_{c}&\Omega^{*0}_{c}\end{array}\right).$ (21)
The masses of the particles involved in the calculation are chosen based on the recommended central values from the Review of Particle Physics (PDG) Tanabashi:2018oca . The mass of the broad $\sigma$ meson is taken as 500 MeV. The coupling constants used in the calculations are listed in Table 1.
Table 1: The coupling constants adopted in the calculation, which are cited from the literature Chen:2019asm ; Liu:2011xc ; Isola:2003fh ; Falk:1992cx . The $\lambda$ and $\lambda_{S,I}$ are in the units of GeV-1. Others are in the units of $1$.
$\beta$ $g$ $g_{V}$ $\lambda$ $g_{s}$ $\ell_{S}$
$0.9$ $0.59$ $5.9$ $0.56$ $0.76$ $6.2$
$\beta_{S}$ $g_{1}$ $\lambda_{S}$ $\beta_{B}$ $\ell_{B}$ $g_{4}$ $\lambda_{I}$
$-1.74$ $-0.94$ $-3.31$ $-\beta_{S}/2$ $-\ell_{S}/2$ $3g_{1}/{(2\sqrt{2})}$ $-\lambda_{S}/\sqrt{8}$
Based on the aforementioned Lagrangians, we can derive the interaction mechanisms responsible for generating molecular strange hidden-charm pentaquarks. To investigate their kaon production dynamics, it is crucial to incorporate couplings between these exotic states and the $\bar{K}N$ channel. While the broad energy region covered here suggests potential contributions from multiple inelastic channels, our current focus remains on providing experimentally testable predictions. Notably, these inelastic channels primarily influence cross-section magnitudes rather than the fundamental production mechanisms. Given the current absence of experimental data for these exotic states, our theoretical predictions target order-of-magnitude estimates - a resolution level sufficient for guiding initial experimental searches. Moreover, the essential characteristics of cross-section behavior (excluding absolute normalization) remain relatively insensitive to these coupling details due to the large mass gap. We therefore employ a simplified one-boson exchange framework to model the crucial coupling between the $\bar{K}N$ channel and charmed-sector final states, maintaining theoretical consistency while preserving predictive capability.
To write the potential for the $\bar{K}N$ interaction, the Lagrangians constructed using heavy quark and chiral symmetries are introduced and presented explicitly as follows Zhu:2022fyb ; Ronchen:2012eg ,
$\mathcal{L}_{\bar{K}\bar{K}{\rho}}$ $=ig_{\bar{K}\bar{K}\rho}\bar{K}^{{\dagger}}{\bm{\tau}}\cdot{\bm{\rho}}^{\mu}\overleftrightarrow{\partial}_{\mu}\bar{K},$
$\mathcal{L}_{\bar{K}\bar{K}{\omega}}$ $=ig_{\bar{K}\bar{K}\omega}\bar{K}^{{\dagger}}{\omega}^{\mu}\overleftrightarrow{\partial}_{\mu}\bar{K},$
$\mathcal{L}_{\bar{K}\bar{K}\sigma}$ $=-g_{\bar{K}\bar{K}\sigma}\bar{K}^{{\dagger}}\sigma\bar{K},$
$\mathcal{L}_{NN\pi}$ $=-\frac{g_{NN\pi}}{m_{\pi}}\bar{N}\gamma^{5}\gamma^{\mu}{\bm{\tau}}\cdot\partial_{\mu}{\bm{\pi}}N,\quad$
$\mathcal{L}_{NN\rho}$ $=-g_{NN\rho}\bar{N}[\gamma^{\mu}-\frac{\kappa_{\rho}}{2m_{N}}\sigma^{\mu\nu}\partial_{\nu}]{\bm{\tau}}\cdot{\bm{\rho}_{\mu}}N,$
$\mathcal{L}_{NN\omega}$ $=-g_{NN\omega}\bar{N}[\gamma^{\mu}-\frac{\kappa_{\omega}}{2m_{N}}\sigma^{\mu\nu}\partial_{\nu}]{\omega_{\mu}}N,$
$\mathcal{L}_{NN\sigma}$ $=-g_{NN\sigma}\bar{N}N\sigma,$ (22)
where $\bar{K}$, $N$, $\pi$, and $\rho$ represent the $\bar{K}$ meson, nucleon, pion meson, and $\rho$ meson fields, respectively. The coupling constants are chosen as follows: $g_{\bar{K}\bar{K}\sigma}=3.65$, $g_{NN\pi}=0.989$, $g_{NN\rho}=-3.1$, $\kappa_{\rho}=1.825$, $\kappa_{\omega}=0$, and $g_{NN\sigma}=5$ Matsuyama:2006rp ; Lu:2020qme .
When calculating the couplings between $\bar{K}N$ and other channels, in addition to the Lagrangians mentioned above, we also need to use the following Lagrangians Wang:2016fhj ; Okubo:1975sc ; Sibirtsev:2000aw ; Liu:2001ce ; Dong:2009tg ; Lu:2016nnt ,
${\cal L}_{DN\Lambda_{c}}$ $=ig_{DN\Lambda_{c}}(\bar{N}\gamma_{5}\Lambda_{c}D+\bar{D}\bar{\Lambda}_{c}\gamma_{5}N),$
${\cal L}_{D^{*}N\Lambda_{c}}$ $=g_{D^{*}N\Lambda_{c}}(\bar{N}\gamma_{\mu}\Lambda_{c}D^{*\mu}+\bar{D}^{*\mu}\bar{\Lambda}_{c}\gamma_{\mu}N),$
${\cal L}_{\Sigma_{c}ND^{*}}$ $=g_{\Sigma_{c}ND^{*}}\bar{N}\gamma_{\mu}{\vec{\tau}}\cdot{\vec{\Sigma}_{c}}D^{*\mu}+\rm{H.c}.,$
${\cal L}_{\Sigma_{c}ND}$ $=-ig_{\Sigma_{c}ND}\bar{N}\gamma_{5}{\vec{\tau}}\cdot{\vec{\Sigma}_{c}}D+\rm{H.c}.$ (23)
The coupling constants, $g_{DN\Lambda_{c}}$ and $g_{D^{N}\Lambda_{c}}$, are chosen as 10.7 and -5.8, respectively Khodjamirian:2011sp . Additionally, $g_{\Sigma_{c}ND^{*}}=3.0$ Dong:2009tg and $g_{\Sigma_{c}ND}=2.69$ Garzon:2015zva .
In the current framework, the inclusion of nine reaction channels naturally leads to 81 pairwise interactions requiring calculation. Explicitly enumerating all these interaction terms would not only prove exceptionally tedious but also introduce significant error-prone risks in formulation. Following the method in Ref. He:2019rva , we input the vertices $\Gamma$ and propagators $P$ directly into the code. The explicit forms of the potential can be written using the Lagrangians and flavor wave functions as follows,
${\cal V}_{\mathbb{P},\sigma}=f_{I}\Gamma_{1}\Gamma_{2}P_{\mathbb{P},\sigma}f(q^{2}),\ \ {\cal V}_{\mathbb{V}}=f_{I}\Gamma_{1\mu}\Gamma_{2\nu}P^{\mu\nu}_{\mathbb{V}}f(q^{2}).$ (24)
The propagators are defined as usual as
$P_{\mathbb{P},\sigma}=\frac{i}{q^{2}-m_{\mathbb{P},\sigma}^{2}},\ \ P^{\mu\nu}_{\mathbb{V}}=i\frac{-g^{\mu\nu}+q^{\mu}q^{\nu}/m^{2}_{\mathbb{V}}}{q^{2}-m_{\mathbb{V}}^{2}}.$ (25)
The form factor $f(q^{2})$ is introduced to account for the off-shell effects of the exchanged meson. It is expressed as $f(q^{2})=e^{-(m_{e}^{2}-q^{2})^{2}/\Lambda_{e}^{2}}$, where $m_{e}$ and $q$ denote the mass and momentum of the exchanged meson, respectively, while $\Lambda_{e}$ serves as a cutoff parameter to mitigate the on-shell effects. In the meson propagator, we substitute $q^{2}\to-|q|^{2}$ to eliminate singularities, following the method outlined in Ref. Gross:2008ps . The factor $f_{I}$ represents the flavor contribution for a given meson exchange in an isoscalar interaction, with the specific values provided in Table 2.
Table 2: The flavor factors $f_{I}$ for certain meson exchanges of certain isoscalar interaction.
$\pi$ $\eta$ $\rho$ $\omega$ $\sigma$
$\bar{K}N\rightarrow\bar{K}N$ $-\frac{3\sqrt{2}}{2}$ $\frac{\sqrt{2}}{2}$ $1$
$\Xi_{c}\bar{D}^{(*)}\rightarrow\Xi_{c}\bar{D}^{(*)}$ $-\frac{3}{2}$ $\frac{1}{2}$ $2$
$\Xi_{c}^{(^{\prime},*)}\bar{D}\rightarrow\Xi_{c}^{(^{\prime},*)}\bar{D}$ $-\frac{3}{4}$ $\frac{1}{4}$ $1$
$\Xi_{c}^{(^{\prime},*)}\bar{D}^{*}\rightarrow\Xi_{c}^{(^{\prime},*)}\bar{D}^{*}$ $-\frac{3}{4}$ $-\frac{1}{12}$ $-\frac{3}{4}$ $\frac{1}{4}$ $1$
$\Lambda_{c}$ $\Sigma_{c}^{(*)}$ $\bar{D}^{(*)}$
$\bar{K}N\rightarrow\Xi_{c}^{(^{\prime},*)}\bar{D^{(*)}}$ $-\frac{\sqrt{2}}{2}$ $-\frac{1}{2}$
$\bar{K}N\rightarrow\Xi_{c}\bar{D}^{(*)}$ $\frac{\sqrt{2}}{2}$
$\bar{K}N\rightarrow\Lambda_{c}\bar{D_{s}}^{(*)}$ $\sqrt{2}$
$D$ $D^{*}$ $D_{s}$ $D_{s}^{*}$
$\bar{D}^{(*)}\Xi_{c}\to J/\psi\Lambda$ $-\sqrt{2}$ $-\sqrt{2}$
$\bar{D}^{(*)}\Xi^{(^{\prime},*)}_{c}\to J/\psi\Lambda$ $-\sqrt{2}$ $-\sqrt{2}$
$\bar{D}_{s}^{(*)}\Lambda_{c}\to J/\psi\Lambda$ $1$ $1$
2.2 The qBSE approach
By inserting the potential kernel into the Bethe-Salpeter equation, the scattering amplitude is derived. The original 4-dimensional integral equation in Minkowski space is simplified to a 1-dimensional integral equation through partial-wave decomposition and the spectator quasipotential approximation, with a fixed spin-parity $J^{P}$, allowing the scattering amplitude to be calculated as He:2015mja ; He:2014nya ; He:2012zd ,
$i{\cal M}^{J^{P}}_{\lambda^{\prime}\lambda}({\rm p}^{\prime},{\rm p})$ $=i{\cal V}^{J^{P}}_{\lambda^{\prime},\lambda}({\rm p}^{\prime},{\rm p})+\sum_{\lambda^{\prime\prime}}\int\frac{{\rm p}^{\prime\prime 2}d{\rm p}^{\prime\prime}}{(2\pi)^{3}}$
$\cdot i{\cal V}^{J^{P}}_{\lambda^{\prime}\lambda^{\prime\prime}}({\rm p}^{\prime},{\rm p}^{\prime\prime})G_{0}({\rm p}^{\prime\prime})i{\cal M}^{J^{P}}_{\lambda^{\prime\prime}\lambda}({\rm p}^{\prime\prime},{\rm p}),\quad\quad$ (26)
where the summation is restricted to nonnegative helicities $\lambda^{\prime\prime}$. The propagator $G_{0}({\rm p}^{\prime\prime})$ is simplified from its original four-dimensional form under the quasipotential approximation and takes the form
$G_{0}$ $=\frac{\delta^{+}(p^{\prime\prime~{}2}_{h}-m_{h}^{2})}{p^{\prime\prime 2}_{l}-m_{l}^{2}}$
$=\frac{\delta^{+}(p^{\prime\prime 0}_{h}-E_{h}({\rm p}^{\prime\prime}))}{2E_{h}({\rm p}^{\prime\prime})[(W-E_{h}({\rm p}^{\prime\prime}))^{2}-E_{l}^{2}({\rm p}^{\prime\prime})]}.$ (27)
As required by the spectator approximation adopted in this work, the heavier particle (denoted as $h$) in a given channel is placed on shell Gross:1999pd , which satisfies $p^{\prime\prime 0}_{h}=E_{h}({\rm p}^{\prime\prime})=\sqrt{m_{h}^{2}+{\rm p}^{\prime\prime 2}}$. The energy of the lighter particle (denoted as $l$) is then given by $p^{\prime\prime 0}_{l}=W-E_{h}({\rm p}^{\prime\prime})$. Here and hereafter, the value of the momentum in center-of-mass frame is defined as ${\rm p}=|{\bm{p}}|$. The partial wave potential is expressed as:
${\cal V}_{\lambda^{\prime}\lambda}^{J^{P}}({\rm p}^{\prime},{\rm p})$ $=2\pi\int d\cos\theta~{}[d^{J}_{\lambda\lambda^{\prime}}(\theta){\cal V}_{\lambda^{\prime}\lambda}({\bm{p}}^{\prime},{\bm{p}})$
$+\eta d^{J}_{-\lambda\lambda^{\prime}}(\theta){\cal V}_{\lambda^{\prime}-\lambda}({\bm{p}}^{\prime},{\bm{p}})],$ (28)
where $\eta=PP_{1}P_{2}(-1)^{J-J_{1}-J_{2}}$, with $P$ and $J$ denoting the parity and spin of the system, as well as those of constituent particles 1 and 2. The relative momenta for the initial and final states are specified as ${\bm{p}}=(0,0,{\rm p})$ and ${\bm{p}}^{\prime}=({\rm p}^{\prime}\sin\theta,0,{\rm p}^{\prime}\cos\theta)$, respectively. The term $d^{J}_{\lambda\lambda^{\prime}}(\theta)$ represents the Wigner d-matrix.
To solve the integral equation (26), the momenta ${\rm p}$, ${\rm p}^{\prime}$, and ${\rm p}^{\prime\prime}$ are discretized using the Gauss quadrature method with weights $w({\rm p}_{i})$. The discretized form of the qBSE can then be expressed as He:2015mja ,
${M}_{ik}$ $={V}_{ik}+\sum_{j=0}^{N}{V}_{ij}G_{j}{M}_{jk}.$ (29)
The propagator $G$ is of a form
$G_{j>0}$ $=\frac{w({\rm p}^{\prime\prime}_{j}){\rm p}^{\prime\prime 2}_{j}}{(2\pi)^{3}}G_{0}({\rm p}^{\prime\prime}_{j}),$
$G_{j=0}$ $=-\frac{i{\rm p}^{\prime\prime}_{\rm o}}{32\pi^{2}W}+\sum_{j}\left[\frac{w({\rm p}_{j})}{(2\pi)^{3}}\frac{{\rm p}^{\prime\prime 2}_{\rm o}}{2W{({\rm p}^{\prime\prime 2}_{j}-{\rm p}^{\prime\prime 2}_{\rm o})}}\right].$ (30)
The ${\rm p}^{\prime\prime}_{\rm o}=\lambda^{1/2}(W,M_{1},M_{2})/2W$ is on-shell momentum with $\lambda(x,y,z)=[x^{2}-(y+z)^{2}][x^{2}-(y-z)^{2}]$ and $W$ represents the total energy of the two-constituent system. To regularize the propagator, an exponential form factor is introduced as $G_{0}({\rm p}^{\prime\prime})\to G_{0}({\rm p}^{\prime\prime})\left[e^{-(p^{\prime\prime 2}_{l}-m_{l}^{2})^{2}/\Lambda_{r}^{4}}\right]^{2}$, where $\Lambda_{r}$ denotes the cutoff He:2015mja . In our framework, the cutoffs applied in the form factors, including those in the propagator and exchange interactions, are treated as free parameters, with $\Lambda_{e}=\Lambda_{r}=\Lambda$. We adopt a cutoff value of $\Lambda=1.04$ GeV, which was successfully used to fit the LHCb data in our previous work Zhu:2022wpi , allowing for slight variations of 0.1 GeV to estimate uncertainties.
The poles of the nine-channel scattering amplitudes for different spin-parity combinations $J^{P}$ are located by varying $z$ in the complex energy plane to satisfy the condition $|1-V(z)G(z)|=0$. For a calculation involving $N$ channels, there are $2^{N}$ Riemann sheets associated with unitarity. In this study, we adopt the method outlined in Ref. Roca:2005nm to identify the poles. Specifically, the default propagator $G(z)$ is used for the energy region below the threshold, while $G(z)+i{\rm p}^{\prime\prime}_{o}/(16\pi^{2}z)$ is adopted for the region above the threshold, corresponding to the first and second Riemann sheets, respectively. This treatment does not affect the physical results on the real energy axis and yields pole positions and half-widths that are closer to those of the corresponding Breit-Wigner forms.
The total cross section is calculated as He:2015cca ,
$\sigma=\frac{1}{16\pi s}\frac{{\rm p}^{\prime}}{\rm p}\sum_{J^{P},\lambda^{\prime}\geq 0\lambda\geq 0}\frac{\tilde{J}}{\tilde{j_{1}}\tilde{j_{2}}}\left|\frac{M^{J^{P}}_{\lambda^{\prime}\lambda}({\rm p}^{\prime},{\rm p})}{4\pi}\right|^{2}.$ (31)
Here, $s$ represents the square of the invariant mass of the initial particle system. The variables $J$ and $j_{i}$ denote the total angular momentum of the system and the spins of the two initial particles, respectively.
3 Numerical Result
3.1 Molecular states without coupled-channel effects
In this section, we study the kaon-induced production of strange hidden-charm pentaquarks, through considering the coupled-channel interactions. However, before conducting a coupled-channel calculation, we first present the results without coupled-channel effects for reference. We analyze the interactions $\bar{K}N$, $\Xi_{c}^{(*)}\bar{D}^{(*)}$, $\Xi^{\prime}_{c}\bar{D}^{(*)}$, $\Lambda_{c}\bar{D}_{s}^{(*)}$, and $\Lambda J/\psi$, while turning off the couplings between these channels. Under this setup, the poles in the complex energy plane correspond directly to the single-channel interaction of each respective channel.
The parameters of the Lagrangians employed here are the same as those used in our previous study on the strange hidden-charm pentaquark states based on LHCb data, specifically $P^{\Lambda}_{\psi s}(4459)$ and $P^{\Lambda}_{\psi s}(4338)$ Zhu:2022wpi . The results are shown in Fig. 1, where the horizontal axis represents the real part of the pole position, and the vertical axis denotes the imaginary part. In this work, only states with spin parity producible from S-wave interactions are considered.
Figure 1: Poles for molecular $P^{\Lambda}_{\psi s}$ state with spin parities $3/2^{-}$ (upper panel) and $1/2^{-}$ (lower panel) obtained without considering coupled-channel effects. The colorboxe represent the values of $\log|1-V(z)G(z)|$. The gray lines correspond to the thresholds of $\Xi^{*}_{c}\bar{D}$, $\Xi_{c}\bar{D}^{*}$, $\Xi^{\prime}_{c}\bar{D}$, $\Lambda_{c}\bar{D}_{s}^{*}$, $\Xi_{c}\bar{D}$, $\Lambda_{c}\bar{D}_{s}$, and $\Lambda J/\psi$, from right to left.
Five poles are identified within the energy range of 4200 to 4550 MeV based on the interactions considered. All poles identified in this work are located below threshold, meaning they reside on the Riemann sheet corresponding to the default propagator $G(z)$ of the respective channel. Since the $\bar{K}N$ threshold is significantly below the charmed energy region and the couplings between channels are turned off, the results closely resemble those from our previous work where the $\bar{K}N$ channel was excluded Zhu:2022wpi . These poles correspond to five molecular states generated from the single-channel interactions of $\Xi_{c}^{*}\bar{D}$ with spin parity $J^{P}=3/2^{-}$, $\Xi_{c}\bar{D}^{*}$ with $1/2^{-}$ and $3/2^{-}$, $\Xi^{\prime}_{c}\bar{D}$ with $1/2^{-}$, and $\Xi_{c}\bar{D}$ with $1/2^{-}$. As the cutoff increases, all poles shift further away from their respective thresholds. Two states with spin parities $1/2^{-}$ and $3/2^{-}$ from the $\Xi_{c}\bar{D}^{*}$ interaction have nearly the same mass as those found in the calculation without coupled-channel effects, approximately 4474 MeV, which is slightly higher than the mass of $P^{\Lambda}_{\psi s}(4459)$. The $\Xi_{c}\bar{D}(1/2^{-})$ interaction generates an S-wave bound state with a mass around 4335 MeV, closely matching the mass of $P^{\Lambda}_{\psi s}(4338)$. The $\Xi^{\prime}_{c}\bar{D}$ interaction produces a molecular state with spin parity $1/2^{-}$. The $\Xi_{c}^{*}\bar{D}$ channel gives rise to a molecular state with spin parity $3/2^{-}$. Even when the cutoff parameter is adjusted to its maximum within a reasonable range, no bound states are formed for the channels $\Lambda_{c}\bar{D}_{s}^{*}$ and $\Lambda_{c}\bar{D}_{s}$.
3.2 Cross section for $J^{P}=1/2^{-}$ in energy region $[4350-4550]$ MeV
In this section, we incorporate the couplings between the interaction channels to study the kaon-induced production of $P^{\Lambda}_{\psi s}$. Since no molecular state is found in the calculations without couplings near the $\Lambda_{c}\bar{D}^{*}$ threshold, the energy regions can clearly be divided into two parts. To present the results more effectively, we provide cross sections for two distinct energy ranges: the higher energy region $[4350-4550]$ MeV and the lower energy region $[4250-4380]$ MeV, where the LHCb observed states, $P^{\Lambda}_{\psi s}(4459)$ and $P^{\Lambda}_{\psi s}(4338)$, are located, respectively. Furthermore, to present results for the molecular pentaquarks with different spin-parities, we focus on the partial wave cross sections rather than the total cross sections.
First, we consider the cross section for $J^{P}=1/2^{-}$ in the energy range $[4350-4550]$ MeV. In the calculation without coupled-channel effects, two poles with spin parity $1/2^{-}$ are generated in this energy region, $\Xi^{\prime}_{c}\bar{D}(1/2^{-})$ and $\Xi_{c}\bar{D}^{*}(1/2^{-})$, with the latter corresponding to $P^{\Lambda}_{\psi s}(4459)$. The imaginary part of the pole with a cutoff of 1.14 GeV is about 10 MeV, corresponding to a width of approximately 20 MeV, which is close to the value reported by LHCb, 17 MeV Aaij:2020gdg . In Fig. 2, after including the coupled-channel effects, the poles for these two states are shown in the complex energy plane. The corresponding cross sections in the channels considered are also presented.
Figure 2: The poles (upper panel) and the cross sections (middle and lower panels) for spin parity $1/2^{-}$, calculated with different cutoffs of 1.04, 1.14, and 1.24 GeV. The gray lines in the upper panel correspond to the thresholds of $\Xi^{*}_{c}\bar{D}$, $\Xi_{c}\bar{D}^{*}$, $\Xi^{\prime}_{c}\bar{D}$, and $\Lambda_{c}\bar{D}_{s}^{*}$, from right to left. In the middle and lower panels, the curves represent the following channels: $\Lambda J/\psi$ (green), $\Xi_{c}\bar{D}$ (blue), $\Xi^{\prime}_{c}\bar{D}$ (black), $\Lambda_{c}\bar{D}_{s}$ (red), and $\Lambda_{c}\bar{D}_{s}^{*}$ (purple).
In the calculation without couplings, two poles are observed on the real axis, positioned just below the thresholds of $\Xi^{\prime}_{c}\bar{D}$ and $\Xi_{c}\bar{D}^{*}$, respectively. When the coupled-channel effects are included, the poles move off the real axis and form two conjugate poles in the complex energy plane. As the parameter increases, the poles shift left along the real axis, indicating stronger binding between the coupled channels. Compared to the results from our previous calculation without the $\bar{K}N$ channel, the pole positions remain similar, as expected, due to the large mass gap between the $\bar{K}N$ channel and the energy region considered here.
The poles in the complex energy plane manifest as structures in the cross section, which can be calculated from the scattering amplitudes along the real axis. As shown in Fig. 2, the cross sections exhibit complex structures due to the coupled-channel interactions in the considered energy region. Generally, the cross sections for the $\Lambda_{c}\bar{D}_{s}$ and $\Lambda_{c}\bar{D}_{s}^{*}$ channels are on the order of tens to hundreds of nanobarns (nb), significantly larger than those for the $\Xi_{c}\bar{D}$ and $\Xi^{\prime}_{c}\bar{D}$ channels, which are on the order of several nb due to different phase spaces and exchange particles as shown in Table 2. The cross section for the $\Lambda J/\psi$ channel is on the order of 1nb or less, which is consistent with predictions from other theoretical studiesClymton:2021thh ; Paryev:2023icm ; Paryev:2022zdx ; Clymton:2022qlr . However, it is significantly smaller than those of other channels, as the $\Lambda J/\psi$ final state cannot couple directly to the $\bar{K}N$ channel.
Although the cross sections for the $\Lambda_{c}\bar{D}_{s}$ and $\Lambda_{c}\bar{D}_{s}^{*}$ channels are large, no obvious structures corresponding to the poles in the complex energy plane (i.e., the molecular pentaquarks) are observed. This suggests that the large cross sections are primarily due to direct scattering rather than the formation of molecular states. As the cutoff $\Lambda$ increases from 1.04 to 1.24 GeV, the cross section for the $\Lambda_{c}\bar{D}_{s}$ channel decreases from about 100 nb to around 10 nb, while the cross section for the $\Lambda_{c}\bar{D}_{s}^{*}$ channel increases from a few nb to about 300 nb. Additionally, a dip is likely to occur around 4450 MeV, attributed to the threshold effect of the $\Xi^{\prime}_{c}\bar{D}$ channel.
For both the $\Xi_{c}\bar{D}$ and $\Xi^{\prime}_{c}\bar{D}$ channels, the cross sections become more prominent. An obvious peak can be seen in the $\Xi^{\prime}_{c}\bar{D}$ channel, with a cross section around 15 nb, which corresponds clearly to the pole below the $\Xi_{c}\bar{D}^{*}$ threshold related to the $P^{\Lambda}_{\psi s}(4459)$ observed at LHCb. However, as the cutoff decreases, the peak diminishes rapidly, and the poles shift towards the $\Xi_{c}\bar{D}^{*}$ threshold. At a cutoff of 1.14 GeV, the peak cross section drops to just a few nb, becoming nearly unobservable at a cutoff of 1.04 GeV. A relatively small peak is also observed in the $\Xi_{c}\bar{D}$ channel at a cutoff of 1.24 GeV, which transitions into a dip at a cutoff of 1.14 GeV, and then a very small peak at a cutoff of 1.04 GeV. This structure originates from the pole below the $\Xi^{\prime}_{c}\bar{D}$ threshold, which also appears as a small peak in the $\Lambda_{c}\bar{D}_{s}^{*}$ channel at cutoffs of 1.24 and 1.14 GeV.
3.3 Cross section for $J^{P}=3/2^{-}$ in energy region $[4350-4550]$ MeV
Now, we consider another partial wave with spin parity $J^{P}=3/2^{-}$ in the same energy region. Without the couplings, two states with $J^{P}=3/2^{-}$ are produced near the $\Xi_{c}\bar{D}^{*}$ and $\Xi^{*}_{c}\bar{D}$ thresholds. After including the coupled-channel effects, the two states move off the real axis and acquire a width, as shown in Fig. 3. Similar to the $1/2^{-}$ case, the cross sections for the channels $\Lambda_{c}\bar{D}_{s}$ and $\Lambda_{c}\bar{D}_{s}^{*}$ are on the order of tens to hundreds of nb, which is much larger than the cross sections for the $\Xi_{c}\bar{D}$ and $\Xi^{\prime}_{c}\bar{D}$ channels, which are on the order of a few nb. The cross section for the $\Lambda J/\psi$ channel is much smaller than that for the other channels.
Figure 3: Similar to Fig. 2, but for the spin-parity $J^{P}=3/2^{-}$.
The higher state near the $\Xi^{*}_{c}\bar{D}$ threshold has a large width. With the increase of the cutoff, the width becomes larger, and the pole in the complex energy plane moves farther from the threshold. A peak corresponding to the pole can be observed in the cross section of the $\Xi^{\prime}_{c}\bar{D}$ channel, and with the increase of the cutoff from 1.04 to 1.24 GeV, the peak becomes higher and more pronounced. However, this pole does not appear in other channels.
The lower state near the $\Xi_{c}\bar{D}^{*}$ threshold, which can be related to the $P^{\Lambda}_{\psi s}(4459)$ observed at LHCb, has a small width. At a cutoff of 1.04 GeV, the pole is close to the real axis and exhibits as a very narrow peak in the cross section of the $\Xi^{\prime}_{c}\bar{D}$ channel. However, with the increase of the cutoff to 1.14 GeV, the pole moves to the $\Xi^{\prime}_{c}\bar{D}$ threshold and naturally can no longer be observed in the $\Xi^{\prime}_{c}\bar{D}$ channel. When the cutoff increases further to 1.24 GeV, the pole will even cross the $\Xi^{\prime}_{c}\bar{D}$ threshold, and the corresponding peak in the $\Xi_{c}\bar{D}$ channel becomes more obvious. This state also exhibits as a small dip in the $\Lambda_{c}\bar{D}_{s}$ channel. Additionally, a dip is also found in the $\Lambda_{c}\bar{D}_{s}$ channel near the $\Xi^{\prime}_{c}\bar{D}$ threshold, as in the case with $1/2^{-}$.
3.4 Cross section for $J^{P}=1/2^{-}$ in energy region $[4250-4380]$ MeV
The calculation without coupling effects suggests the presence of a bound state originating from the $\Xi_{c}\bar{D}$ interaction, corresponding to $P^{\Lambda}_{\psi s}(4338)$. As shown in Fig.4, after including coupled-channel effects, the state acquires a small imaginary part corresponding to a width of about 5 MeV, consistent with the LHCb observation LHCb:2022ogu . When the cutoff parameter $\Lambda$ increases, the pole shifts leftward along the real axis, and the pole near the $\Xi_{c}\bar{D}$ threshold becomes more prominent.
Figure 4: Similar to Fig. 2, but in a lower energy region.
Below $4338$ MeV, there are only two channels, the $\Lambda J/\psi$ and the $\Lambda_{c}\bar{D}_{s}$, that we need to consider. From the calculation results of the scattering cross section, regardless of the value of the parameter, the scattering cross section in the $\Lambda J/\psi$ channel is very small and almost undetectable. The scattering cross section results in the $\Lambda_{c}\bar{D}_{s}$ channel show a swing structure occurring around $4340$ MeV at all cuoffs of 1.04, 1.14 and 1.24 GeV. This is due to the pole near the $\Xi_{c}\bar{D}$ threshold, which corresponds to $P^{\Lambda}_{\psi s}(4338)$, with interference from the background of other interactions.
4 Summary
In this study, we investigated the kaon-induced production of strange hidden-charm molecular pentaquark states, including $P^{\Lambda}_{\psi s}(4338)$ and $P^{\Lambda}_{\psi s}(4459)$, using a coupled-channel calculation. The interactions considered involved the $\bar{K}N$, $\Xi_{c}^{(*)}\bar{D}^{(*)}$, $\Xi^{\prime}_{c}\bar{D}^{(*)}$, $\Lambda_{c}\bar{D}_{s}^{(*)}$, and $\Lambda J/\psi$ channels. Effective Lagrangians were employed to construct the interaction potentials, which were incorporated into the quasipotential Bethe-Salpeter equation to compute scattering amplitudes. Partial wave cross sections for $J^{P}=1/2^{-}$ and $3/2^{-}$ are analyzed, revealing complex structures and significant energy-dependent behavior.
For partial wave $J^{P}=1/2^{-}$, in the lower energy region, the pole near the $\Xi_{c}\bar{D}$ threshold, associated with $P^{\Lambda}_{\psi s}(4338)$, generates a noticeable swing structure in the cross section for the $\Lambda_{c}\bar{D}_{s}$ and $\Lambda_{c}\bar{D}_{s}^{*}$ channels. This structure remains stable across variations in the cutoff and exhibits a magnitude of several nb. In the higher energy region, the $\Lambda_{c}\bar{D}_{s}$ and $\Lambda_{c}\bar{D}_{s}^{*}$ channels exhibit large cross sections, reaching up to several hundreds of nb. However, no distinct peaks corresponding to molecular states were observed, apart from threshold-induced cusps, suggesting that these contributions arise primarily from direct scattering processes rather than resonance effects.
In the $\Xi^{\prime}_{c}\bar{D}$ channel, a prominent peak, associated with a pole below the $\Xi_{c}\bar{D}^{*}$ threshold, is observed with a cross section of approximately 15 nb. This peak is linked to the $P^{\Lambda}_{\psi s}(4459)$ state observed by LHCb. As the cutoff decreases, the peak diminishes significantly, becoming barely visible. In the $\Xi_{c}\bar{D}$ channel, a smaller peak is evident at a cutoff of 1.24 GeV, transitions into a dip at 1.14 GeV, and reappears as a very weak peak at 1.04 GeV. These structures originate from the pole below the $\Xi^{\prime}_{c}\bar{D}$ threshold, which also gives rise to small peaks in the $\Lambda_{c}\bar{D}_{s}^{*}$ channel at higher cutoff values.
For the partial wave $J^{P}=3/2^{-}$, the pole near the $\Xi_{c}\bar{D}^{*}$ threshold, associated with the $P^{\Lambda}_{\psi s}(4459)$, manifests as a peak in the $\Xi^{\prime}_{c}\bar{D}$ and $\Xi_{c}\bar{D}$ channels. At a cutoff of 1.04 GeV, this pole exhibits a narrow width, remaining close to the real axis, and produces a sharp peak in the $\Xi^{\prime}_{c}\bar{D}$ channel. As the cutoff increases, the pole shifts to the $\Xi^{\prime}_{c}\bar{D}$ threshold and disappears from the $\Xi^{\prime}_{c}\bar{D}$ channel. At even higher cutoffs, the pole crosses the $\Xi^{\prime}_{c}\bar{D}$ threshold, resulting in a narrow peak in the $\Xi_{c}\bar{D}$ channel. This state also generates a small dip in the $\Lambda_{c}\bar{D}_{s}$ channel. Additionally, states near the $\Xi_{c}^{*}\bar{D}$ threshold produce a broad peak in the $\Xi^{\prime}_{c}\bar{D}$ channel, characterized by a large width and relatively weak sensitivity to the cutoff. Furthermore, a dip near the $\Xi^{\prime}_{c}\bar{D}$ threshold is observed in the $\Lambda_{c}\bar{D}_{s}$ channel, similar to the behavior seen in the $J^{P}=1/2^{-}$ case. These observations highlight the intricate interplay between thresholds and poles, driven by coupled-channel dynamics.
These findings suggest that the strange hidden-charmed pentaquarks, such as $P^{\Lambda}_{\psi s}(4338)$ and $P^{\Lambda}_{\psi s}(4459)$, may potentially be observed in kaon-induced production processes. There are still many alternative assignments for the $P^{\Lambda}_{\psi s}$ Wang:2021itn ; Peng:2020hql ; Feijoo:2022rxf . Studying hidden-charmed pentaquarks through new production mechanisms, from both theoretical and experimental perspectives, may also be beneficial for distinguishing these different assignments. High-precision experiments utilizing kaon beams at J-PARC and JLab would be instrumental in confirming the existence of these states and exploring their properties in detail.
Acknowledgement J. He is supported by the National Science Foundation of China (Grant No. 12475080), and J. T. Zhu is supported by the Start-up Funds of Changzhou University (Grant No. ZMF24020043) and the National Science Foundation of China (Grant No. 12405090).
Data Availability Statement This manuscript has no associated data or the data will not be deposited. [Authors’ comment: This is a theoretical study and no external data are associated with this work.]